equation of focal chord of parabola y^2=4x inclined at an angle 45 with x axis is?????

Grade:12

1 Answers

Arun
5 years ago
Dear student
 
focus of y^2 = 4x is (1,0)
 
focal chord inclined at an angle of 45 degree from x axis
 
y – 0 = 1(x- 1)
 
y = x – 1
 
Hope it helps
 
Regards
Arun (askIITians forum expert)